# Thomason condition for regular algebraic stacks

### v2

@@ 4. Proofs
This section proves our results. To start, we need a few lemmas.
@@
Let $f\colon\mathcal{Y}\to\mathcal{X}$ be a morphism of quasi-compact quasi-separated algebraic stacks. Consider a closed subset $Z\subseteq|\mathcal{X}|$. Then
@@
Hence, we can reduce to the case of affine schemes. In this setting, $D_{\operatorname{qc},Z}(\mathcal{X})$ is compactly generated (see e.g. [Rou08, Theorem 6.8]). From Remark˜3.2, we know that $D_{\operatorname{qc},Z}(\mathcal{X})$ is compactly generated by a subset $\mathcal{B}$ of $D_{\operatorname{qc},Z}(\mathcal{X})\cap\operatorname{Perf}(\mathcal{X})$. Furthermore, every object of $D_{\operatorname{qc},Z}(\mathcal{X})$ is a homotopy colimit of iterated extensions of small coproducts of shifts of objects in $\mathcal{B}$ (see [Sta25, Tag 09SN]). So, the desired claim can be checked if we can show $\mathbf{L}f^{\ast}\mathcal{B}\subseteq D_{\operatorname{qc},f^{-1}(Z)}(\mathcal{Y})$. However, [HR17a, Lemma 4.8(2)] tells us
which completes the proof. ∎
Lemma 4.2.
Let $F\colon\mathcal{T}\leftrightarrows\mathcal{S}\colon G$ be an adjoint pair of exact functors between triangulated categories admitting small coproducts. Assume $\mathcal{T}$ is compactly generated by a collection $\mathcal{B}$ and $G$ commutes with small coproducts. Then $G$ is conservative (i.e. $G(A)\cong 0\implies A\cong 0$) if, and only if, $F(\mathcal{B})$ compactly generates $\mathcal{S}$. In such a case, if $\mathcal{T}$ is compactly generated by a set of cardinality $\leq\beta$ for some cardinal $\beta$, then $\mathcal{B}$ satisfies the same condition.
Proof.
This is known to some but we spell it out for convenience. That $G$ commutes with small coproducts ensures that $F(\mathcal{T}^{c})\subseteq\mathcal{S}^{c}$ (see e.g. the proof of $\implies$ in [Nee96, Theorem 5.1]; which this fact does not require $\mathcal{T}$ to be compactly generated). First, assume $G$ is conservative. Let $E\in\mathcal{B}$ satisfy $\operatorname{Hom}(F(B),E[n])\cong 0$ for all $B\in\mathcal{B}$ and $n\in\mathbb{Z}$. From adjunction, it follows that $\operatorname{Hom}(B,G(E)[n])\cong 0$. As $\mathcal{B}$ compactly generates $\mathcal{T}$, it follows that $G(E)\cong 0$. However, $G$ being conservative implies $E\cong 0$. So, $F(\mathcal{B})$ compactly generates $\mathcal{S}$.
Conversely, assume that $F(\mathcal{B})$ compactly generates $\mathcal{S}$. Let $E\in\mathcal{S}$ such that $G(E)\cong 0$. By adjunction, it follows that $0\cong\operatorname{Hom}(F(B),E[n])$ for all $B\in\mathcal{B}$ and $n\in\mathbb{Z}$. Yet, the assumption implies $E\cong 0$. Hence, $G$ must be conservative.
That the last claim holds follows from the proof above. ∎
Lemma 4.3.
Let $t\colon\mathcal{Y}\to\mathcal{X}$ be an immersion of quasi-compact quasi-separated algebraic stacks. If $\mathcal{X}$ satisfies the $\beta$-Thomason condition, then so does $\mathcal{Y}$.
Proof.
This is [Nee23, Observation 5.6], but we give a slightly different proof. An immersion is quasi-affine, and hence, concentrated. So, [HR17a, Lemma 8.2] guarantees that $D_{\operatorname{qc}}(\mathcal{Y})$ is compactly generated by a collection of at most $\leq\beta$ objects. Here, $t$ is injective on the underlying topological spaces. Let $Z$ be a closed subset of $|\mathcal{Y}|$. Set $Z^{\prime}$ to be the closure of $t(Z)$ in $|\mathcal{X}|$. Then $t^{-1}(Z^{\prime})=Z$. Moreover, the hypothesis on $\mathcal{X}$ allows us to find a $P\in\operatorname{Perf}(\mathcal{X})$ such that $\operatorname{supp}(P)=Z^{\prime}$. However, [HR17a, Lemma 4.8(2)] tells us $t^{-1}(\operatorname{supp}(P))=\operatorname{supp}(\mathbf{L}t^{\ast}P)$, which completes the proof because $\mathbf{L}t^{\ast}P\in\operatorname{Perf}(\mathcal{Y})$. ∎
Lemma 4.4.
Let $f\colon\mathcal{Y}\to\mathcal{X}$ be a finite morphism of Noetherian algebraic stacks. If $D_{\operatorname{qc}}(\mathcal{Y})$ is compactly generated by $\mathcal{B}\subseteq D_{\operatorname{qc}}(\mathcal{Y})^{c}$, then the adjoint pair $\mathbf{R}f_{\ast}$ and $f^{!}$ restricts to $D_{\operatorname{qc}}(\mathcal{Y})$ and $D_{\operatorname{qc},f(|\mathcal{Y}|)}(\mathcal{X})$. Additionally, if $\mathbf{R}f_{\ast}\mathcal{O}_{\mathcal{Y}}\in\operatorname{Perf}(\mathcal{X})$, then $\mathbf{R}f_{\ast}\mathcal{B}$ compactly generates $D_{\operatorname{qc},f(|\mathcal{Y}|)}(\mathcal{X})$.
Proof.
@@
To start, we show the claim regarding restricting the adjunction. Recall, by [HR17a, Theorem 4.14(1)], the right adjoint of $\mathbf{R}f_{\ast}$ on $D_{\operatorname{qc}}$ exists because $f$ is concentrated. It is easy to see that $f^{!}(D_{\operatorname{qc},f(|\mathcal{Y}|)}(\mathcal{X}))\subseteq D_{\operatorname{qc}}(\mathcal{Y})$. So, we only need to show that $\mathbf{R}f_{\ast}(D_{\operatorname{qc}}(\mathcal{Y}))\subseteq D_{\operatorname{qc},f(|\mathcal{Y}|)}(\mathcal{X})$. Let $E\in D_{\operatorname{qc}}(\mathcal{Y})$. Since $D_{\operatorname{qc}}(\mathcal{Y})$ is compactly generated by $\mathcal{B}$, [Kra22, Corollary 3.4.9] (or [Sta25, Tag 09SN]) tells us $E$ can be written as the homotopy colimit of a sequence
$E_{0}\xrightarrow{\phi_{0}}E_{1}\xrightarrow{\phi_{1}}E_{2}\xrightarrow{\phi_{2}}\cdots$
such that $E_{0}\cong 0$ and $\operatorname{cone}(\phi_{i})$ is a small coproduct of shifts of objects in $\mathcal{B}$. Now, as $\mathbf{R}f_{\ast}$ preserves homotopy colimits, it suffices to show $\mathbf{R}f_{\ast}E_{i}$ and $\operatorname{cone}(\mathbf{R}f_{\ast}\phi_{i})$ belong to $D_{\operatorname{qc},f(|\mathcal{Y}|)}(\mathcal{X})$. Using an inductive argument on $i$, it suffices to show small coproducts of shifts of objects from $\mathcal{B}$ satisfy the same condition. Furthermore, $\mathbf{R}f_{\ast}$ preserving small coproducts and shifts, so we can reduce to checking the same condition on just the objects of $\mathcal{B}$. However, this can be seen by reducing to schemes (see e.g. [GW23, Remark 23.46(2)]), e.g. $\mathbf{R}f_{\ast}\mathcal{B}\subseteq D^{b}_{\operatorname{coh},f(|\mathcal{Y}|)}(\mathcal{X})$.
@@
Lastly, we show that $\mathbf{R}f_{\ast}\mathcal{B}$ compactly generates $D_{\operatorname{qc},f(|\mathcal{Y}|)}(\mathcal{X})$. Since [HR17a, Theorem 4.14(4)] tells us $f^{!}$ preserves small coproducts, Lemma˜4.2 ensures that it suffices to show that $f^{!}\colon D_{\operatorname{qc},f(|\mathcal{Y}|)}(\mathcal{X})\to D_{\operatorname{qc}}(\mathcal{Y})$ is conservative. So, let $E\in D_{\operatorname{qc},f(|\mathcal{Y}|)}(\mathcal{X})$ satisfy $f^{!}E\cong 0$. By [HR17a, Theorem 4.14(2)], it follows that
$0\cong\mathbf{R}f_{\ast}f^{!}E\cong\operatorname{\mathbf{R}\mathcal{H}\!\mathit{om}}(\mathbf{R}f_{\ast}\mathcal{O}_{\mathcal{Y}},E).$
@@
However, $\mathbf{R}f_{\ast}\mathcal{O}_{\mathcal{Y}}\in\operatorname{Perf}(\mathcal{X})\cap D_{\operatorname{qc},f(|\mathcal{Y}|)}(\mathcal{X})$, and so, [HR17a, Lemma 4.9] ensures that $E\cong 0$. Hence, we see that the restriction of $f^{!}$ on $D_{\operatorname{qc},f(|\mathcal{Y}|)}(\mathcal{X})$ is conservative. ∎
@@
Recall that a collection $\mathcal{B}$ in a category $\mathcal{C}$ is said to generate $\mathcal{C}$ if for any nonzero $E\in\mathcal{C}$ there is an $n\in\mathbb{Z}$ and $B\in\mathcal{B}$ such that $\operatorname{Hom}(B[n],E)\not\cong 0$.
Theorem 4.6.
Let $\mathcal{X}$ be a quasi-compact quasi-separated regular algebraic stack with quasi-finite and locally separated diagonal. Then there is an object of $\operatorname{Perf}(\mathcal{X})$ which generates $D_{\operatorname{qc}}(\mathcal{X})$.
Proof.
Using [Ryd11, Theorem 7.2], we can find an étale, surjective, finitely presented (i.e. of finite presentation), representable by algebraic spaces morphism $p\colon\mathcal{Y}\to\mathcal{X}$ and a finite, flat, surjective, finitely presented morphism $v\colon U\to\mathcal{Y}$ from a quasi-affine scheme. Furthermore, appealing to [Ryd11, Proposition 4.4], there is a finite sequence of quasi-compact open immersions
such that the base change of $f$ to $|\mathcal{X}_{c}|\setminus|\mathcal{X}_{c-1}|$ (when given any closed substack structure) is étale, finite, and of constant rank. Note that every algebraic stack above is Noetherian.
By inducting on $n$, we can reduce to the setting where $f$ is finite after base change along a closed immersion $t\colon\mathcal{Z}\to\mathcal{X}$ and the open substack $\mathcal{U}:=|\mathcal{X}|\setminus t(|\mathcal{Z}|)$ satisfies $D_{\operatorname{qc}}(\mathcal{U})$ being generated by an object $G_{U}\in\operatorname{Perf}(\mathcal{U})$. Set $Z=t(|\mathcal{Z}|)$ and let $j\colon\mathcal{U}\to\mathcal{X}$ be the associated open immersion. Now, consider the commutative diagram obtained by base changes,
[…]
where each internal square is fibered. Here, $f^{\prime}$ and $v^{\prime}$ are flat, finite, and surjective morphisms. Since $t^{\prime\prime}$ is closed immersion, we have that $U\times_{\mathcal{X}}\mathcal{Z}$ is a Noetherian scheme. So, there is a $P\in\operatorname{Perf}(U\times_{\mathcal{X}}\mathcal{Z})$ which compactly generates $D_{\operatorname{qc}}(U\times_{\mathcal{X}}\mathcal{Z})$. Then Lemma˜4.4 (or Proposition˜A.1) tells us that $\mathbf{R}(f^{\prime}\circ v^{\prime})_{\ast}P$ compactly generates $D_{\operatorname{qc}}(\mathcal{Z})$ as $f^{\prime}\circ v^{\prime}$ is flat, finite, and surjective.
By [HLLP25, Proposition B.1], there is a Verdier localization sequence
$D^{b}_{\operatorname{coh},Z}(\mathcal{X})\xrightarrow{i_{\ast}}D^{b}_{\operatorname{coh}}(\mathcal{X})\xrightarrow{\mathbf{L}j^{\ast}}D^{b}_{\operatorname{coh}}(\mathcal{U}).$
However, [DLMP25, Theorem 3.7] tells us that $\operatorname{Perf}=D^{b}_{\operatorname{coh}}$ in each case. This allows us to find a $G\in\operatorname{Perf}(\mathcal{X})$ such that $\mathbf{L}j^{\ast}G\cong G_{U}$; that is, $\mathbf{L}j^{\ast}G$ generates $D_{\operatorname{qc}}(\mathcal{U})$ and belongs to $\operatorname{Perf}(\mathcal{U})$. We claim that $G\oplus\mathbf{R}(f^{\prime}\circ v^{\prime})_{\ast}P$ generates $D_{\operatorname{qc}}(\mathcal{X})$. Towards this end, consider the recollement obtained in Proposition˜3.1,
$D_{\operatorname{qc}}(\mathcal{U})$$D_{\operatorname{qc}}(\mathcal{X})$$D_{\operatorname{qc},Z}(\mathcal{X})$$\scriptstyle{\mathbf{R}j_{\ast}}$$\scriptstyle{\mathbf{L}j^{\ast}}$$\scriptstyle{j^{!}}$$\scriptstyle{i^{!}}$$\scriptstyle{i_{\ast}}$$\scriptstyle{i^{\ast}}$
where in particular $i_{\ast}$ is the natural inclusion and $j^{!}$ the right adjoint of $\mathbf{R}j_{\ast}$. Let $E\in D_{\operatorname{qc}}(\mathcal{X})$ satisfy $\operatorname{Hom}(G\oplus\mathbf{R}(f^{\prime}\circ v^{\prime})_{\ast}P[n],E)=0$ for all $n\in\mathbb{Z}$. Since $i_{\ast}$ is the inclusion, we have symbolically that $\mathbf{R}(f^{\prime}\circ v^{\prime})_{\ast}P=i_{\ast}\mathbf{R}(f^{\prime}\circ v^{\prime})_{\ast}P$. From the recollement above, there is a distinguished triangle
$i_{\ast}i^{!}E\to E\to\mathbf{R}j_{\ast}\mathbf{L}j^{\ast}E\to i_{\ast}i^{!}E[1].$
Here, in such a case, we have that $\operatorname{Hom}(i_{\ast}\mathbf{R}(f^{\prime}\circ v^{\prime})_{\ast}P[n],E)\cong 0$ for all $n\in\mathbb{Z}$. So, adjunction tells us $\operatorname{Hom}(\mathbf{R}(f^{\prime}\circ v^{\prime})_{\ast}P[n],i^{!}E)\cong 0$ for all $n\in\mathbb{Z}$. However, $\mathbf{R}(f^{\prime}\circ v^{\prime})_{\ast}P$ generates $D_{\operatorname{qc},Z}(\mathcal{X})$, so $i^{!}E\cong 0$. This tells us the morphism $E\to\mathbf{R}j_{\ast}\mathbf{L}j^{\ast}E$ from the distinguished triangle above is an isomorphism. Once more, from adjunction, we then have that $\operatorname{Hom}(G[n],\mathbf{R}j_{\ast}\mathbf{L}j^{\ast}E)=0$ for all $n\in\mathbb{Z}$. So, $\operatorname{Hom}(\mathbf{L}j^{\ast}G[n],\mathbf{L}j^{\ast}E)=0$ via adjunction for any such $n$. However, $\mathbf{L}j^{\ast}G$ generates $D_{\operatorname{qc}}(\mathcal{U})$, so $\mathbf{L}j^{\ast}E\cong 0$. Consequently, $E\cong 0$, which completes the proof. ∎
Corollary 4.7.
Let $\mathcal{X}$ be a concentrated regular algebraic stack with quasi-finite and locally separated diagonal. Then $\mathcal{X}$ satisfies the $1$-Thomason condition. Moreover, if $\mathcal{S}\to\mathcal{X}$ is an immersion, then $\mathcal{S}$ satisfies the $1$-Thomason condition.
Proof.
That the last claim holds if we can prove the first by Lemma˜4.3. As $\mathcal{X}$ is regular and Noetherian, it follows that $\operatorname{Perf}(\mathcal{X})=D^{b}_{\operatorname{coh}}(\mathcal{X})$ (see e.g. [DLMP25, Theorem 3.7]). Let $W\subseteq|\mathcal{X}|$ be closed. Denote by $i\colon\mathcal{W}\to\mathcal{X}$ for the closed immersion from the reduced induced closed substack structure on $W$. If $\mathcal{X}$ is concentrated, then $\mathbf{R}i_{\ast}\mathcal{O}_{\mathcal{W}}\in D_{\operatorname{qc}}(\mathcal{X})^{c}$ and has support coinciding with $W$. Moreover, by Theorem˜4.6, there is an $E\in D^{b}_{\operatorname{coh}}(\mathcal{X})$ which generates $D_{\operatorname{qc}}(\mathcal{X})$. If $\mathcal{X}$ is concentrated, then $\operatorname{Perf}(\mathcal{X})=D_{\operatorname{qc}}(\mathcal{X})^{c}$, and so $D_{\operatorname{qc}}(\mathcal{X})$ is compactly generated as desired. ∎
Appendix A Descending Thomason condition
While not needed in proving our main results, we record the following for sake of interest (cf. [HR17a, Theorem 6.6] in the case for $\beta$-crispness).
Proposition A.1.
@@ Appendix A Descending Thomason condition
Since $f\colon\mathcal{Y}\to\mathcal{X}$ is a finite, flat, surjective morphism of Noetherian algebraic stacks, it follows that $\mathbf{R}f_{\ast}\mathcal{O}_{\mathcal{Y}}\in\operatorname{Perf}(\mathcal{X})$. Denote by $f^{!}$ the right adjoint of $\mathbf{R}f_{\ast}$ on $D_{\operatorname{qc}}$. In our case, [HR17a, Theorem 4.14] tells us that $f^{!}$ preserves small coproducts and is conservative. By the hypothesis, we know that there exists a collection $\mathcal{B}$ (of some cardinality $\leq\beta$) which compactly generates $D_{\operatorname{qc}}(\mathcal{Y})$.
@@
We claim that the collection of $\mathbf{R}f_{\ast}(B\otimes^{\mathbf{L}}P)$ with $B\in\mathcal{B}$ compactly generates $D_{\operatorname{qc},Z}(\mathcal{X})$. Towards that end, we show that the adjunction $\mathbf{R}f_{\ast}\dashv f^{!}$ restricts to an adjoint pair between $D_{\operatorname{qc},f^{-1}(Z)}(\mathcal{Y})$ and $D_{\operatorname{qc},Z}(\mathcal{X})$. By [HR17a, Theorem 4.14(4)], we know that $f^{!}\mathcal{O}_{\mathcal{X}}\otimes^{\mathbf{L}}\mathbf{L}f^{\ast}E\cong f^{!}E$ for all $E\in D_{\operatorname{qc},Z}(\mathcal{X})$. This ensures that $\operatorname{supp}(f^{!}(E))\subseteq\operatorname{supp}(\mathbf{L}f^{\ast}E)$. Moreover, Lemma˜4.1 promises that $\operatorname{supp}(\mathbf{L}f^{\ast}E)\subseteq f^{-1}(Z)$, and so $f^{!}D_{\operatorname{qc},Z}(\mathcal{X})\subseteq D_{\operatorname{qc},f^{-1}(Z)}(\mathcal{Y})$.

### v3

@@ 3.2. A special case
The following is recorded for interest sake.
Corollary 3.3.
Let $\mathcal{X}$ be a quasi-compact quasi-separated algebraic stack. Suppose $j\colon\mathcal{U}\to\mathcal{X}$ is a quasi-compact open immersion. Set $Z:=|\mathcal{X}|\setminus|\mathcal{U}|$. For any closed $W\subseteq|\mathcal{X}|$ such that $D_{\operatorname{qc},|\mathcal{U}|\cap W}(\mathcal{U})$ is compactly generated, there exists a recollement
$D_{\operatorname{qc},|\mathcal{U}|\cap W}(\mathcal{U})$$D_{\operatorname{qc},W}(\mathcal{X})$$D_{\operatorname{qc},W\cap Z}(\mathcal{X})$$\scriptstyle{\mathbf{R}j_{\ast}}$$\scriptstyle{\mathbf{L}j^{\ast}}$$\scriptstyle{j^{!}}$$\scriptstyle{i^{!}}$$\scriptstyle{i_{\ast}}$$\scriptstyle{i^{\ast}}$
where $i_{\ast}$ is the natural inclusion.
Proof.
This mimics the proof of Proposition˜3.1 but a few details need to be spelled out. By Lemma˜2.1, the restriction of $\mathbf{L}j^{\ast}$ to $D_{\operatorname{qc},W}(\mathcal{X})$ factors through $D_{\operatorname{qc},|\mathcal{U}|\cap W}(\mathcal{U})$. Moreover, the restriction of $\mathbf{R}j_{\ast}$ to $D_{\operatorname{qc},|\mathcal{U}|\cap W}(\mathcal{U})$ factors through $D_{\operatorname{qc},W}(\mathcal{X})$, which can be verified using [HR17a, Theorem 2.6(2)] and checking smooth locally. Hence, the adjoint pair $\mathbf{L}j^{\ast}$ and $\mathbf{R}j_{\ast}$ restricts to $D_{\operatorname{qc},W}(\mathcal{X})$ and $D_{\operatorname{qc},|\mathcal{U}|\cap W}(\mathcal{U})$. Furthermore, there is a Verdier localization
$D_{\operatorname{qc},W\cap Z}(\mathcal{X})\xrightarrow{i_{\ast}}D_{\operatorname{qc},W}(\mathcal{X})\xrightarrow{\mathbf{L}j^{\ast}}D_{\operatorname{qc},|\mathcal{U}|\cap W}(\mathcal{U}).$
Indeed, the counit of $\mathbf{L}j^{\ast}$ and $\mathbf{R}j_{\ast}$ is a natural isomorphism on $D_{\operatorname{qc}}$, and so the restriction $\mathbf{L}j^{\ast}\colon D_{\operatorname{qc},W}(\mathcal{X})\to D_{\operatorname{qc},|\mathcal{U}|\cap W}(\mathcal{U})$ is essentially surjective. To see that fully faithfulness holds, use e.g. [GZ67, §I. Proposition 1.3] and [KRA10, Lemma 4.3.1]). Note that restriction of $\mathbf{R}j_{\ast}$ is right adjoint to the restriction of $\mathbf{L}j^{\ast}$. So, by [GP18, Lemma 2.2(ii)], $i_{\ast}$ must admit a right adjoint as well, which we denote by $i^{!}$ (loc. cit. pulls from [CPS88b, Theorem 1.1] and [CPS88a, Theorem 2.1]). It follows that
$D_{\operatorname{qc},Z,\cap W}(\mathcal{X})\xleftarrow{i^{!}}D_{\operatorname{qc},W}(\mathcal{X})\xleftarrow{\mathbf{R}j_{\ast}}D_{\operatorname{qc},|\mathcal{U}|\cap W}(\mathcal{U})$
is a Verdier localization sequence. However, $D_{\operatorname{qc},|\mathcal{U}|\cap W}(\mathcal{U})$ is compactly generated and the restriction of $\mathbf{R}j_{\ast}$ preserves small coproducts. Hence, [BDS16, Corollary 2.3] tells us the restriction of $\mathbf{R}j_{\ast}$ admits a right adjoint. Now, we can apply [GP18, Lemma 2.2(ii)] once more to see that $i^{!}$ admits a right adjoint as well, which we denote by $i^{\ast}$. Tying things together, we have the required data for a recollement; i.e. a Verdier localization sequence which is a localization and colocalization sequence in the sense of [KRA10, §4]. ∎
Remark 3.4.
In Corollary˜3.3, the restriction of $\mathbf{R}j_{\ast}$ to $D_{\operatorname{qc},|\mathcal{U}|\cap W}(\mathcal{U})\to D_{\operatorname{qc},W}(\mathcal{X})$ admits a right adjoint. However, it is not clear to us whether it is the restriction of the right adjoint $j^{\times}$ of $\mathbf{R}j_{\ast}$ on $D_{\operatorname{qc}}$.
3.3. Generating
We record a few useful lemmas allowing one to glue generators along a recollement. Recall that a collection $\mathcal{B}$ in a category $\mathcal{C}$ is said to generate $\mathcal{C}$ if for any nonzero $E\in\mathcal{C}$ there is an $n\in\mathbb{Z}$ and $B\in\mathcal{B}$ such that $\operatorname{Hom}(B[n],E)\not\cong 0$.
Lemma 3.5.
Let $F\colon\mathcal{T}\leftrightarrows\mathcal{S}\colon G$ be an adjoint pair of exact functors between triangulated categories admitting small coproducts. Suppose $\mathcal{T}$ is generated by a collection $\mathcal{B}$. Then $G$ is conservative (i.e. $G(A)\cong 0\implies A\cong 0$) if, and only if, $F(\mathcal{B})$ generates $\mathcal{S}$. In such a case, if $\mathcal{T}$ is generated by a set of cardinality $\leq\beta$ for some cardinal $\beta$, then $\mathcal{B}$ satisfies the same condition.
Proof.
This is known to a few but we spell it out for convienence. Let $E\in\mathcal{B}$ satisfy $\operatorname{Hom}(F(B),E[n])\cong 0$ for all $B\in\mathcal{B}$ and $n\in\mathbb{Z}$. From adjunction, we know that $\operatorname{Hom}(B,G(E)[n])\cong 0$. As $\mathcal{B}$ generates $\mathcal{T}$, it follows that $G(E)\cong 0$. However, $G$ being conservative implies $E\cong 0$. So, $F(\mathcal{B})$ generates $\mathcal{S}$.
Conversely, assume that $F(\mathcal{B})$ generates $\mathcal{S}$. Let $E\in\mathcal{S}$ satisfy $G(E)\cong 0$. By adjunction, it follows that $0\cong\operatorname{Hom}(F(B),E[n])$ for all $B\in\mathcal{B}$ and $n\in\mathbb{Z}$. Yet, the assumption implies $E\cong 0$. Hence, $G$ must be conservative.
That the last claim holds follows from the proof above. ∎
Corollary 3.6.
Let $F\colon\mathcal{T}\leftrightarrows\mathcal{S}\colon G$ be an adjoint pair of exact functors between triangulated categories admitting small coproducts. Suppose $\mathcal{T}$ is compactly generated by a collection $\mathcal{B}$ and $G$ commutes with small coproducts. Then $G$ is conservative if, and only if, $F(\mathcal{B})$ compactly generates $\mathcal{S}$. In such a case, if $\mathcal{T}$ is compactly generated by a set of cardinality $\leq\beta$ for some cardinal $\beta$, then $\mathcal{B}$ satisfies the same condition.
Proof.
That $G$ commutes with small coproducts ensures that $F(\mathcal{T}^{c})\subseteq\mathcal{S}^{c}$ (see e.g. the proof of $\implies$ in [NEE96, Theorem 5.1]; which this fact does not require $\mathcal{T}$ to be compactly generated). Consequently, the desired claim follows from Lemma˜3.5. ∎
Proposition 3.7.
Consider a recollement as in Diagram˜3.1. Let $\mathcal{G}$ be a subcategory of $\mathcal{K}$ such that $I_{\lambda}(\mathcal{G})$ generates $\mathcal{T}$. If $\mathcal{G}^{\prime}$ is a subcategory which generates $\mathcal{D}$, then the collection of $G\oplus Q_{\lambda}(G^{\prime})$ where $G\in\mathcal{G}$ and $G^{\prime}\in\mathcal{G}^{\prime}$ generates $\mathcal{K}$.
Proof.
Let $E\in\mathcal{K}$ satisfy $\operatorname{Hom}((G\oplus Q_{\lambda}(G^{\prime}))[n],E)=0$ for all $n\in\mathbb{Z}$, $G\in\mathcal{G}$, and $G^{\prime}\in\mathcal{G}^{\prime}$. From the recollement above, there is a distinguished triangle
$(Q_{\lambda}\circ Q)(E)\to E\to(I\circ I_{\lambda})(E)\to(Q_{\lambda}\circ Q)(E)[1].$
Here, in such a case, we have that $\operatorname{Hom}(Q_{\lambda}(G^{\prime})[n],E)\cong 0$ for all $n\in\mathbb{Z}$. So, adjunction tells us $\operatorname{Hom}(G^{\prime}[n],Q(E))\cong 0$ for all $n\in\mathbb{Z}$. However, $\mathcal{G}^{\prime}$ generates $\mathcal{D}$, so $Q(E)\cong 0$. This implies the morphism $E\to(I\circ I_{\lambda})(E)$ from the distinguished triangle above is an isomorphism. Once more, from adjunction, we then have that $\operatorname{Hom}(G[n],(I\circ I_{\lambda})(E))=0$ for all $n\in\mathbb{Z}$. So, $\operatorname{Hom}(I_{\lambda}(G)[n],I_{\lambda}(E))=0$ via adjunction for any such $n$. However, $I_{\lambda}(\mathcal{G})$ generates $\mathcal{T}$, so $I_{\lambda}(E)$. Consequently, $E\cong 0$, which completes the proof. ∎
@@ 4. Proofs
This section containes our main results. To start, we prove the following. It is useful in later proofs where we leverage a duality trick.
@@
Let $f\colon\mathcal{Y}\to\mathcal{X}$ be a finite morphism of Noetherian algebraic stacks. Then the adjoint pair $\mathbf{R}f_{\ast}$ and $f^{\times}$ restricts to $D_{\operatorname{qc}}(\mathcal{Y})$ and $D_{\operatorname{qc},f(|\mathcal{Y}|)}(\mathcal{X})$. Additionally, if $D_{\operatorname{qc}}(\mathcal{Y})$ is generated by some $\mathcal{B}\subseteq D^{b}_{\operatorname{coh}}(\mathcal{X})$ and $\mathbf{R}f_{\ast}\mathcal{O}_{\mathcal{Y}}\in\operatorname{Perf}(\mathcal{X})$, then $\mathbf{R}f_{\ast}\mathcal{B}$ generates $D_{\operatorname{qc},f(|\mathcal{Y}|)}(\mathcal{X})$.
@@
To start, we show the claim regarding restricting the adjunction. Note that a finite morphism is a closed mapping on the underlying topological spaces. Recall, by [HR17a, Theorem 4.14(1)], the right adjoint of $\mathbf{R}f_{\ast}$ on $D_{\operatorname{qc}}$ exists because $f$ is concentrated. It is easy to see that $f^{\times}(D_{\operatorname{qc},f(|\mathcal{Y}|)}(\mathcal{X}))\subseteq D_{\operatorname{qc}}(\mathcal{Y})$. Also, by Lemma˜2.2, we know that $\mathbf{R}f_{\ast}(D_{\operatorname{qc}}(\mathcal{Y}))\subseteq D_{\operatorname{qc},f(|\mathcal{Y}|)}(\mathcal{X})$ because $f$ is finite.
@@
Lastly, we show that $\mathbf{R}f_{\ast}\mathcal{B}$ generates $D_{\operatorname{qc},f(|\mathcal{Y}|)}(\mathcal{X})$. By Lemma˜3.5, it suffices to show that $f^{\times}\colon D_{\operatorname{qc},f(|\mathcal{Y}|)}(\mathcal{X})\to D_{\operatorname{qc}}(\mathcal{Y})$ is conservative. So, let $E\in D_{\operatorname{qc},f(|\mathcal{Y}|)}(\mathcal{X})$ satisfy $f^{\times}E\cong 0$. By [HR17a, Theorem 4.14(2)], it follows that
$0\cong\mathbf{R}f_{\ast}f^{\times}E\cong\operatorname{\mathbf{R}\mathcal{H}\!\mathit{om}}(\mathbf{R}f_{\ast}\mathcal{O}_{\mathcal{Y}},E).$
@@
However, $\mathbf{R}f_{\ast}\mathcal{O}_{\mathcal{Y}}\in\operatorname{Perf}(\mathcal{X})\cap D_{\operatorname{qc},f(|\mathcal{Y}|)}(\mathcal{X})$, and so, [HR17a, Lemma 4.9] ensures that $E\cong 0$. Hence, we see that the restriction of $f^{\times}$ on $D_{\operatorname{qc},f(|\mathcal{Y}|)}(\mathcal{X})$ is conservative. ∎
@@
In Lemma˜4.1, [HR17a, Theorem 4.14(4)] tells us $f^{\times}$ preserves small coproducts. So, Corollary˜3.6 tells us $\mathbf{R}f_{\ast}\mathcal{B}$ compactly generates $D_{\operatorname{qc},f(|\mathcal{Y}|)}(\mathcal{X})$ if $\mathcal{B}$ does such for $D_{\operatorname{qc}}(\mathcal{Y})$.
Next, we prove a variation of [HR17a, Theorem 6.6] for the $\beta$-Thomason condition. Although it is used in the proof of Theorem˜1.1, it is possible to replace references to it with [HR17a, Theorem 6.6]. So, we record it for interest sake.
Proposition 4.3.
@@
Since $f\colon\mathcal{Y}\to\mathcal{X}$ is a finite, flat, surjective morphism of Noetherian algebraic stacks, it follows that $\mathbf{R}f_{\ast}\mathcal{O}_{\mathcal{Y}}\in\operatorname{Perf}(\mathcal{X})$. Denote by $f^{\times}$ the right adjoint of $\mathbf{R}f_{\ast}$ on $D_{\operatorname{qc}}$. In our case, [HR17a, Theorem 4.14] tells us that $f^{\times}$ preserves small coproducts and is conservative. By the hypothesis, we know that there exists a collection $\mathcal{B}$ (of some cardinality $\leq\beta$) which compactly generates $D_{\operatorname{qc}}(\mathcal{Y})$.
@@
We claim that the collection of $\mathbf{R}f_{\ast}(B\otimes^{\mathbf{L}}P)$ with $B\in\mathcal{B}$ compactly generates $D_{\operatorname{qc},Z}(\mathcal{X})$. Towards that end, we show that the adjunction $\mathbf{R}f_{\ast}\dashv f^{\times}$ restricts to an adjoint pair between $D_{\operatorname{qc},f^{-1}(Z)}(\mathcal{Y})$ and $D_{\operatorname{qc},Z}(\mathcal{X})$. By [HR17a, Theorem 4.14(4)], we know that $f^{\times}\mathcal{O}_{\mathcal{X}}\otimes^{\mathbf{L}}\mathbf{L}f^{\ast}E\cong f^{\times}E$ for all $E\in D_{\operatorname{qc},Z}(\mathcal{X})$. This ensures that $\operatorname{supp}(f^{\times}(E))\subseteq\operatorname{supp}(\mathbf{L}f^{\ast}E)$. Moreover, Lemma˜2.1 promises that $\operatorname{supp}(\mathbf{L}f^{\ast}E)\subseteq f^{-1}(Z)$, and so $f^{\times}D_{\operatorname{qc},Z}(\mathcal{X})\subseteq D_{\operatorname{qc},f^{-1}(Z)}(\mathcal{Y})$.